How many 5 letter code words can be formed using a 27 letter alphabet?
A-repetitions allowed
B-no repetitions allowed

Answer by Theo(13342)

if repetitions are allowed, then 27^5 = 14348907
if repetitions are not allowed, then 27P5 = 27! / 22! = 9687600

how's this work?

in order to see, try 2 letter code words from 3 letter alphabet of abc.

with repetition = 3^2 = 9
without repetition = 3P2 = 3! / 1! = 6

with repetition, the codes would be:

aa
ab
ac
ba
bb
ac
ca
cb
cc

total of 9.

without repetition, the codes would be:

ab
ac
ba
bc
ca
cb

the same formulas apply to the bigger problem.

the formula for permutation is nPx = n! / (n-x)!